OpenStudy (anonymous):

simplify the following radical: 3√72 √ = radical sign

OpenStudy (blank ):

18 sqrt(2)

OpenStudy (blank ):

3 sqrt(72) Pull all perfect square roots out from under the radical. In this case, remove the 6 because it is a perfect square. (3*6 sqrt(2)) Multiply 3 by 6 to get 18. 18 sqrt(2) Hope this helps :)

OpenStudy (anonymous):

\[3\sqrt{72}\\\\3\sqrt{36\times2}=2\sqrt{36}\sqrt{2}\\3\times6\sqrt2\\\\\boxed{18\sqrt2}\]

OpenStudy (anonymous):

In this case, 72 can be divided by 2 into 36, which is a perfect square. So: \[3\sqrt{72}=3\sqrt{36\times2}\] Which then equals \[3\sqrt{36}\sqrt{2}\] So then, you multiply 3 with the perfect square of 36, which is 6. \[3\times6\sqrt2\] Then, simplified, equals \[18\sqrt2\] And that's how I got my answer *bows*